Find the area under normal curve from z = -2.83 to z= -0.82. 


1
Expert's answer
2021-05-07T11:57:45-0400

"P(-2.83<Z<-0.82)=P(Z<-0.82)-P(Z<-2.83)="

"=P(Z<2.83)-P(Z<0.82)=0.9977-0.7939=0.2038."